OpenStudy (anonymous):

11x^6/8-^2 * 64p^5/121x multiply and simplify. x in the second equation belongs with 121 only.

OpenStudy (anonymous):

\[11x^6/8p^2 * 64p^5/121x\] that is how the problem should look.

OpenStudy (anonymous):

\[\frac{11x^6}{8p^2}*\frac{64p^5}{121x}\] \[=\frac{8x^5p^3}{11}\]

OpenStudy (anonymous):

hmm my assignment said that answer is incorrect.

OpenStudy (anonymous):

is that p^2 or p^3 ?

OpenStudy (anonymous):

ok I got it. thanks